Let x be a random variable with the following probability distribution Value x of x P(X=x -10 0.05 0 0.20 10 0.30 20 0.20 30 0.10 40 0.15 E (x)= Var (x)=
Answer #1

The PMF table is given below.

X -10 0 10 20 30 40
P(X) 0.05 0.2 0.3 0.2 0.10 0.15

The expected value is

E(X) = xP(X = x) E(X) = -10(0.05)+0(0.2)+10(0.3)+20(0.2)+30(0.10)+40(0.15) E(X) = 15.5

To find variance,

E(X2) = x?P(X = x) E(X) = 102(0.05) + 0(0.2) + 10P(0.3) + 20P(0.2) + 302(0.10) + 402(0.15) E(X) = 445 Var(x) = E(X2) - E(X)?